Paragraph Styles On November 1, 2013 John Adams, a customer service representative of Americo Drilling Supplies (ADS), was summoned to the Houston office of Drilling Contractors, Inc. (DCI), to inspect three boxcars of mud-treating agents that ADS had shipped to the Houston firm. DCI had filed a complaint that the 25 Pound bags of treating agents that it had just received from ADS were light-weighted. The light-weight bags were initially detected by one of DCIs receiving clerks, who noticed that the rail road side scale tickets indicated that the net weights were significantly less on all three of the boxcars than those of identical shipments received on October 25, 2013. ADSs traffic department was called to determine if lighter-weight dunnage or pallets were used on the shipments. (This might explain the lighter weights.) ADS indicated, however, that no changes had been made in the loading or palletizing procedures. Hence, DCI randomly checked 25 of the bags and discovered that the average net weight was almost 24.5 pounds. Consequently, they concluded that the sample indicated a significance short-weight. contacted, and Adams was sent to investigate the complaint and to issue to DCI. ADS, was then credit (?2?) DCI, however, was not completely satisfied with only the issuance of credit for the short shipment t. The charts followed by their mud engineers on the drilling platforms 46 144 8 4 5 6 8 9
media%2F10c%2F10cff2a0-b2c0-4f88-88fd-a1
media%2F522%2F52239766-a5f8-4534-87fd-c8
media%2Fbdb%2Fbdb12d2a-c91e-436e-a8ad-19
media%2Fdd7%2Fdd7c8b3c-d7e9-4364-bfab-5d
media%2F203%2F2034f62e-2815-4b5d-8aaf-52
media%2Fd11%2Fd1151b01-9b20-4018-9fc9-f6
media%2F792%2F7927b935-2970-44a0-a32c-85
media%2F926%2F926d11f9-1701-41c3-bc81-57
media%2F210%2F21077aa1-190c-4814-b20f-bd
media%2F58f%2F58fe7834-a31a-475d-8155-ee
Answer #1

(a)ans.

To find out the values for range column by using following formula:

Range = Maximum – Minimum = Largest – Smallest

Calculate the range of column; we have to find the means and standard deviations for the given 4 columns such as the average weights, smallest weights, largest weights, range weights.

Required means and standard deviations by using excel are summarized as below:

(Range column is calculated by using Excel operations.)

Average weight

Smallest

Largest

Range

Sample size

72

72

72

72

Total

1766.8

1706.3

1830

123.7

Mean

24.53888889

23.69861111

25.41666667

1.718055556

Standard deviation

0.586840563

0.831677933

0.7218911

0.741396928

(b)ans.

σι-bar- Vn Oz-bar-standard deviation of x-bar σ = standard deviation of the population n = sample-size-10 Oz-bar = 0.5868 σ-0.5868 * V10 1.8558 So, standard devlatlon of Indivldual bags (1.e. of population of Indlvidual bags) 1.8558

(d)ans

The control limits are determined as follows:

Central line = CL = overall mean = Xbar = 24.54 and range mean = Rbar = 1.72

For sample size of n = 10, A2 = 0.31, D3 = 0.22 and D4 = 1.78

UCLXbar = Xbar + (A2)(Rbar) = 24.54 + (0.31)(1.72) = 25.07

LCLXbar = Xbar - (A2)(Rbar) = 24.54 - (0.31)(1.72) = 24.006

UCLXbar = 25.07 and LCLXbar = 24.006

UCLR = (D4)(Rbar) = (1.78)(1.72) = 3.058

LCLXbar = (D3)(Rbar) = (0.22)(1.72) = 0.38

UCLR = 3.058 and LCLR = 0.38

(e)ans

Compute Xbarbar (Average of Bar of 72 samples) and Rebar (Average of Range of 72 samples)

Rest of the steps of statistical process control is following


Conclusion:

3 R values are outside the control limits. Therefore, based on R-chart, the process is not in control.

36 Xbar values are outside the control limits. Therefore, based on Xbar-chart, the process is not in control